OpenStudy (anonymous):

E is the midpoint of DF, DE=2x+4, and EF = 3x-1. Find DE, EF, and DF

OpenStudy (abmon98):

3x-1=2x+4 3x-2x=4+1 x=5

OpenStudy (abmon98):

DE=2x+4-->2(5)+4=14 EF=3x-1--->3(5)-1=14 DF is EF added to DE 14+14=28